To whom may it concern,
I have this tricky question to solve.
If x + 2y = -2 and x – 2y = 18, find the value of x^2 + 4y^2.
According to my teacher, the answer is 164 but how ?

Please help. Thanks.
CH Lam

Answer by tutorcecilia(2152)

Remember that y=y and x=x, than
Solve for y (or x) in each equation:
x + 2y = -2 [solve for y]
y=-1/2x-1
.
x – 2y = 18 [solve for y]
y=1/2x-9
.
Since y=y, than
-1/2x-1=1/2x-9 [Solve for x]
1x=8
x=8
.
Plugging (x=8) back into either of the original equations:
x + 2y = -2
(8)+2y=-2 [Solve for y]
y=-10/2
y=-5
.
.
Plug both values (x=8) and (y=-5) into the problem:
x^2+4y^2
(8)^2+4(-5)^2
64+4(25)
64+100
164